Circular No. 6694

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1997 N1 (TABUR)

The following precise CCD positions have been reported:

1997 UT α_{2000} δ_{2000} m_{1} Observer
July 4.33101 4^{h}59^{m}27.^{s}00 −26^{o}29′44′′.3 12.5 Holvorcem
4.33852 4 59 30.32 −26 29 34.0 12.7  _{”}
5.79969 5 10 18.94 −25 58 19.0  Garradd
5.82001 5 10 27.95 −25 57 51.6  _{”}

P. R. Holvorcem (Barão Geraldo). 0.20-m f/6.3 reflector. G. J. Garradd (Loomberah). 0.25-m f/4.1 reflector.

Parabolic orbital elements from 14 observations July 3–5:

T = 1997 Aug. 15.247 TT  ω = 345.^{o}448
Ω = 148.240 }2000.0
q = 0.38322 AU  i = 86.402
1997TT α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{1}
July 1 4^{h}35.^{m}46 −27^{o}27.^{′}9 1.256 1.136 58.^{o}9 50.^{o}0 10.0
6 5 11.80 −25 53.8 1.202 1.043 55.3 53.3  9.6
11 5 49.69 −23 36.7 1.166 0.947 50.9 56.4  9.1
16 6 27.84 −20 34.3 1.150 0.850 45.7 58.8  8.6
21 7 04.95 −16 49.8 1.156 0.752 39.8 60.0  8.1
26 7 40.05 −12 29.5 1.182 0.654 33.6 59.1  7.5
31 8 12.64 − 7 39.2 1.225 0.559 26.9 55.1  6.9

SUPERNOVA 1997cu IN ANONYMOUS GALAXY The Mount Stromlo Abell Cluster Supernova Search Team (cf. IAUC

6639) reports the discovery of an apparent supernova (V = 18.2, R = 18.0) on CCD images taken on July 4 by S. Sabine at the Mount Stromlo 1.27-m telescope (+ Macho Camera). SN 1997cu is located near Abell 3128 at α = 3^{h}29^{m}03^{s}.0, δ = −52^{o}41′41′′ (equinox 2000.0), within 1′′ of its highsurface-brightness host. SN 1997cu was not present on similar frames taken on Feb. 25 to R = 20. A spectrum (range 360–750 nm) obtained on the MSSSO 2.3-m reflector + DBS spectrograph on July 5 reveals that SN 1997cu is of type Ia and near maximum. The redshift of the host galaxy, measured from several absorption lines is z = 0.062, coincident with the redshift of Abell 3128 (z = 0.06).

1997 July 5  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Brian G. Marsden